# Thermodynamic topological classes of the rotating, accelerating black holes

### v1

@@ 5 Concluding remarks
In this paper, we extend our prior work PRD108-084041 by employing the generalized off-shell Helmholtz free energy to explore the topological characteristics of rotating, accelerating black holes. Specifically, we analyze the topological number of the Kerr C-metric and Kerr-AdS C-metric black holes. Additionally, we apply the same approach to examine the topological number of the KN C-metric and KN-AdS C-metric black holes. Our findings reveal that the Kerr C-metric and KN C-metric black holes share the same topological classification, both having a topological number of $W=1$, while the Kerr-AdS C-metric and KN-AdS C-metric black holes belong to a different topological class, characterized by a topological number of $W=0$.
Building on the results presented in this paper and our previous study PRD108-084041 , we identify three intriguing outcomes:
  1. 1.
    The topological number of an asymptotically flat accelerating black hole consistently differs by one from that of its non-accelerating counterpart.
  2. 2.
    For an asymptotically AdS accelerating black hole, the topological number is reduced by one compared to the corresponding non-accelerating AdS black hole.
  3. 3.
    In the context of general relativity, the acceleration parameter and the negative cosmological constant each independently raise the topological number by one. However, when both are present, their influences counteract, leaving the topological number unchanged.
There are two promising directions for future research emerge from this work. The first is to investigate the topological numbers of black holes in alternative theories of gravity, including scalar-tensor-vector gravity Moffat:2005si ; Liu:2023uft ; 2406.00579 ; Qiao:2024gfb , bumblebee gravity Casana:2017jkc ; Liu:2022dcn ; Liu:2024oeq , and Kalb-Ramond gravity Yang:2023wtu ; 2406.13461 ; 2407.07416 , etc. We have noted that Wei et al. 2409.09333 recently proposed a universal thermodynamic topological classification method for black hole solutions. Therefore, the second outlook is to explore the universal thermodynamic topological classes of rotating and accelerating black holes.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12205243, No. 12375053, No. 12475051, No. 12122504, and No. 12035005; the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347; the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028; the science and technology innovation Program of Hunan Province under Grant No. 2024RC1050; the innovative research group of Hunan Province under Grant No. 2024JJ1006; and the Hunan Provincial Major Sci-Tech Program under Grant No.2023ZJ1010.

### v2

@@ V Concluding remarks
In this paper, by employing the generalized off-shell Helmholtz free energy, we explore the topological characteristics of rotating, accelerating black holes. Specifically, we analyze the topological number of the Kerr C-metric and Kerr-AdS C-metric black holes. Additionally, we apply the same approach to examine the topological number of the KN C-metric and KN-AdS C-metric black holes. Our findings reveal that the Kerr C-metric and KN C-metric black holes share the same topological classification, both having a topological number of $W=1$, while the Kerr-AdS C-metric and KN-AdS C-metric black holes belong to a different topological class, characterized by a topological number of $W=0$.
Building on the results presented in this paper and our previous study PRD108-084041 , we identify three intriguing outcomes: I) The topological number of an asymptotically flat accelerating black hole consistently differs by one from that of its non-accelerating counterpart. II) For an asymptotically AdS accelerating black hole, the topological number is reduced by one compared to the corresponding non-accelerating AdS black hole. III) In the context of general relativity, the acceleration parameter and the negative cosmological constant each independently raise the topological number by one. However, when both are present, their influences counteract, leaving the topological number unchanged.
In summary, by studying the thermodynamic topological classification of rotating and accelerating black holes, we find that their thermodynamic phase transition behaviors can be uniformly characterized through topological invariants (e.g., winding numbers, topological numbers), revealing the regulatory mechanisms of angular momentum and acceleration on their thermodynamic properties. By transforming phase transition problems into the analysis of structures in topological space, this classification not only systematizes our understanding of known phase transitions (such as the Hawking-Page transition and P-V criticality) but also predicts novel thermodynamic phase transition behaviors 2402.00106 ; 2405.07525 , bridging the connection between macroscopic thermodynamics and microscopic quantum structures. This approach is increasingly becoming a crucial tool for exploring the nature of black holes and gravity theories.
There are two promising directions for future research emerge from this work. The first is to investigate the topological numbers of black holes in alternative theories of gravity, including scalar-tensor-vector gravity Moffat:2005si ; Liu:2023uft ; 2406.00579 ; Qiao:2024gfb , bumblebee gravity Casana:2017jkc ; Liu:2022dcn ; Liu:2024oeq , and Kalb-Ramond gravity Yang:2023wtu ; 2406.13461 ; 2407.07416 (additional fields may introduce new topological defects, potentially altering the value of $W$). We have noted that Refs. 2409.09333 ; 2409.12747 ; 2411.10102 ; 2504.10858 recently proposed a universal thermodynamic topological classification method for black hole solutions. Therefore, the second outlook is to explore the universal thermodynamic topological classes of rotating and accelerating black holes.
@@ Acknowledgements.
We are greatly indebted to the anonymous referees for their constructive comments to improve the presentation of this work. This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12205243, No. 12375053, No. 12475051, No. 12122504, and No. 12035005; the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347; the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028; the science and technology innovation Program of Hunan Province under Grant No. 2024RC1050; the innovative research group of Hunan Province under Grant No. 2024JJ1006; and the Hunan Provincial Major Sci-Tech Program under Grant No.2023ZJ1010.